OpenStudy (anonymous):

FAN/MEDAL. Find an equation with integral coefficients that has the given roots: 1, -3, -4

OpenStudy (anonymous):

a: x3 + 12x2 + 5x - 12 = 0 b: x3 + 6x2 + 5x - 12 = 0 c: x3 - 6x2 + 5x - 12 = 0 d: x3 - 12x2 + 10x - 24 = 0

OpenStudy (anonymous):

so you write it as (x-1)(x+3)(x+4)

OpenStudy (anonymous):

solve it you will get the answer

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

welcome :)